OpenStudy (anonymous):

Please I really need help with my Algebra 2 homework

OpenStudy (anonymous):

(x, y) = 1

OpenStudy (anonymous):

whats the question?

OpenStudy (anonymous):

Okay let me send it to you

OpenStudy (anonymous):

Find three ordered pairs that satisfy the equation. y = 3x − 3

OpenStudy (anonymous):

(x, y) = 1

OpenStudy (anonymous):

I need the answer to this question

OpenStudy (anonymous):

Find three ordered pairs that satisfy the equation y=3x-3 First, put 1 in for x and solve. y=3(1)-3 y=3-3 y=0 (1, 0) Next, put in a 2 for x and solve. y=3(2)-3 y=6-3 y=3 (2,3) Lastly, put in a 3 for x and solve. y=3(3)-3 y=9-3 y=6 (3,6) This gives you three ordered pairs. They are (1,0), (2,3), and (3,6). I hope this helped.

OpenStudy (anonymous):

Thanks but could please stay with me and help me because it really hard to find someone to help me

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

but for answer x,y=

OpenStudy (anonymous):

(1,

OpenStudy (anonymous):

which one was it because am a bit confused on what you send me

OpenStudy (anonymous):

is the answer for x,y=(1,

OpenStudy (anonymous):

I am little bit confused for that part. because if its asking for coordinates then its (1,0)

OpenStudy (anonymous):

I see do think the answer maybe (1,0

OpenStudy (anonymous):

oh just 0

OpenStudy (anonymous):

(x, y) = 1

OpenStudy (anonymous):

I am asking for little help here. @GreenCat can you help?

OpenStudy (anonymous):

@uri

OpenStudy (anonymous):

am good now it was 0

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

for the next one is this Find three ordered pairs that satisfy the equation. y = x + 4

OpenStudy (anonymous):

(x, y) = 0, (x, y) = 1, (x, y) = 2

OpenStudy (anonymous):

you can do it same way as we did above. just put values of x = 0,1,2 and so on to find y

OpenStudy (anonymous):

Went you say values do I multiply everything together that's were am confused am not good in math especially algebra

OpenStudy (anonymous):

Find three ordered pairs that satisfy the equation y=x+4 First, put 0 in for x and solve. y=0+4 y=4 (0, 4) Next, put in a 1 for x and solve. y=1+4 y=5 (1,5) Lastly, put in a 2 for x and solve. y=2+4 y=6 (2,6)

OpenStudy (anonymous):

Okay thanks let me try

OpenStudy (anonymous):

Okay are you still there

OpenStudy (anonymous):

yeah